OpenStudy (anonymous):

is the equation always sometimes or never true 15+2x-4=9x+11-7x 2x+3(4x-1)=2(5x+3)+4x

OpenStudy (anonymous):

a. never true b. always true B. a. always true b. sometimes true C. a. always true b. never true D. a. sometimes true b. never true

OpenStudy (anonymous):

a. never true b. always true B. a. always true b. sometimes true C. a. always true b. never true D. a. sometimes true b. never true

OpenStudy (anonymous):

15+2x-4=9x+11-7x 15-4+2x=9x-7x+11 11+2x=11+2x

OpenStudy (anonymous):

15+2x-4 = 9x+11-7x Combine the like terms

OpenStudy (anonymous):

Which is always true right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x+3(4x-1)=2(5x+3)+4x 2x+12x-3=10x+6+4x 14x-3=14x+6

OpenStudy (anonymous):

So what do u think about this

OpenStudy (anonymous):

always true

OpenStudy (anonymous):

I dont think so

OpenStudy (anonymous):

so are they suppose to equal each other , SOMETIMES ?

OpenStudy (lgbasallote):